OpenStudy (anonymous):

logb(sqrt(27))-logb(sqrt(32))

OpenStudy (anonymous):

what is the goal here? are you supposed to write this as a single log?

OpenStudy (anonymous):

I'll attach question

OpenStudy (anonymous):

or are you supposed to expand it even further? for example \[\sqrt{27}=3\sqrt3\] and \[\sqrt{32}=4\sqrt2\]

OpenStudy (anonymous):

yeah that would be good, see the actual question

OpenStudy (anonymous):

well that is a whole different story

OpenStudy (anonymous):

since \[\sqrt{27}=3\sqrt3\] we have \[\log(\sqrt{27})=\log(3\sqrt3)=\log(3)+\log(\sqrt3)=\log(3)+\frac{1}{2}\log(3)\]

OpenStudy (anonymous):

since you are told that \(\log(3)=C\) then this translates to \[c+\frac{1}{2}C=\frac{3}{2}C\]

OpenStudy (anonymous):

similar idea for \[\log(\sqrt{32}=\log(4)+\frac{1}{2}\log(2)\]

OpenStudy (anonymous):

oh damn i made a mistake, it is \(\log(3)=A\) not \(C\) no matter, change the first thing to \[\frac{3}{2}A\]

OpenStudy (anonymous):

then the second part is \[\log(4)+\frac{1}{2}\log(2)=2\log(2)+\frac{1}{2}\log(2)=2C+\frac{1}{2}C=\frac{5}{2}C\]

OpenStudy (anonymous):

that was exhausting final answer \[\frac{3}{2}A-\frac{5}{2}C\]

OpenStudy (anonymous):

thank you so much. can i ask you about another problem too? because wolfram and i got the same answer but its wrong in webworks so i dont understand